Cisplatin or Platinol has a chemical PtCl2(NH3)2 that is used in the treatment of certain cancers. Calculate (a) the amount (mol) of compound in 285.3 g of cisplatin; (b) the number of hydrogen atoms in 0.98 mol of cisplatin.


1
Expert's answer
2022-03-17T14:58:28-0400

M(cisplatin) = M(PtCl2(NH3)2) = 300 g/mol;

If m(cisplatin) = 285.3 g

Then:

n(PtCl2(NH3)2) = m(PtCl2(NH3)2)/M(PtCl2(NH3)2) = 285.3/300 = 0.951 moles;

Na = 6.02 * 1023 mol-1;

If n(PtCl2(NH3)2) = 0.98 mol

Then:

n(H) = 6 * n(PtCl2(NH3)2) = 5.88 mol;

N(H) = n(H) * Na = 5.88 * 6.02 * 1023 = 35.4 * 1023 = 3.54 * 1024.

Answer: (a) 0.951 moles;

(b) 3.54 * 1024.